A particle is observed to have a lifetime of 1.50 X 10^-6 s when it is at rest in the laboratory. What is its lifetime when it is moving at 0.800c? How far will the particle move with respect to the moving frame of reference before it decays? How far will the particle move with respect to the laboratory frame before it decays?


1
Expert's answer
2020-09-29T09:39:36-0400

Solution

Lifetime in moving frame is given by

"t'=\frac{t}{\sqrt{1-\frac{v^2}{c^2}}}"


"t'=\frac{1.50\times10^{-6}}{\sqrt{1-\frac{(0.800c)^2}{c^2}}}"

"t'=2.5\times 10^{-6} s"

distance travelled by particle before decays is

"d=v\times t'=0.800c\times2.5\times10^{-6} \\ d=600 m"